OpenStudy (anonymous):

In the voting for city council precinct 5, only 1/2 of all eligible voters cast votes. what fraction of all eligible voters voted for shelley ? daley ? who received the most votes? ~Will earn medal

OpenStudy (anonymous):



OpenStudy (anonymous):

please answer

OpenStudy (mathmale):

Let x be the TOTAL number of eligible voters. Only half of these eligible voters actually vote, so the number of actual voters is (x/2). Shelley gets (1/10)th of the votes of the actual voters, or (1/10)*(x/2). Simplify this.

OpenStudy (anonymous):

I don't understand

OpenStudy (mathmale):

Can't read y our table. Copy your original table and paste it. Tell me, please, what you DO understand, and we'll work from that point onward.

OpenStudy (anonymous):



OpenStudy (mathmale):

Candidate Fraction of votes rec'd Shelley 1/10 Daley 2/8

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Again: Let x be the number of eligible voters (not all of them vote!) Since only half of the eligible voters vote, the number who vote is (x/2). Daley receives 2/8 of the votes actually cast. In other words, Daley receives (2/8)*(x/2) votes. \[\frac{ 2 }{ 8 }*\frac{ x }{ 2 }=\frac{ 2x }{ 8 }\]

OpenStudy (mathmale):

Please reduce this last result.

OpenStudy (anonymous):

1/4

OpenStudy (mathmale):

Actually, that'd be x/4, or (1/4)x. This is the number of votes that Daley actually receives. How many votes does the other guy receive? (1/10)*(x/2) = ??

OpenStudy (anonymous):

the question says how many VOTERS were there for daley

OpenStudy (anonymous):

hello

OpenStudy (mathmale):

We found the number of voters who voted for Daley just recently. Earlier, we found the number who voted for the other person. How many voters voted for each person? Type them into your table.

OpenStudy (mathmale):

Candidate Fraction of votes rec'd Actual number of votes received Shelley 1/10 Daley 2/8

OpenStudy (anonymous):

the question just has that data table

OpenStudy (mathmale):

All the answers you need are in our conversation; please go back and look for them. Thanks.

OpenStudy (mathmale):

Let x be the TOTAL number of eligible voters. Only half of these eligible voters actually vote, so the number of actual voters is (x/2). Shelley gets (1/10)th of the votes of the actual voters, or (1/10)*(x/2).

OpenStudy (mathmale):

How many votes does Daley get?

OpenStudy (anonymous):

2/8

OpenStudy (anonymous):

I don't understand '-'

OpenStudy (mathmale):

Daley gets 2/8 of the total number of votes. The total number of votes is x/2. Multiply these two fractions together to get the number of votes rec'd by Daley.

OpenStudy (mathmale):

If you don't understand, please ask questions, so that I may help you get back on track.

OpenStudy (anonymous):

you have to find how many people voted for shelly and daley

OpenStudy (mathmale):

Yes, and we have done that twice for each person. See above. What is the total number of votes actually cast? I wrote that out for you several times.

OpenStudy (anonymous):

the answer is shelley : 1/10 dalley : 1/40

OpenStudy (mathmale):

I've explained several times that the total number of votes ACTUALLY CAST is x/2, where x is the number of eligible voters: 1) The total number of votes is x/2. 2) Let x be the TOTAL number of eligible voters. Only half of these eligible voters actually vote, so the number of actual voters is (x/2). 3) Since only half of the eligible voters vote, the number who vote is (x/2). If Daley gets 2/8 of the total number of votes, he gets \[\frac{ 2 }{ 8 }*\frac{ x }{ 2 },~or \frac{ x }{ 4 },\] votes. I am sorry you find this difficult and frustrating. i do ask you to stick with this, as well as to ask questions if I say things that you don't understand.

OpenStudy (anonymous):

I am in 5th grade

OpenStudy (mathmale):

I understand. As before, I'm sorry this topic has been difficult for you. Still, I believe you can understand it. We are using Algebra, and normally, Algebra is taught in the 8th or 9th grades.

OpenStudy (anonymous):

I got 5 voters for shelley and 1/8 for daley

OpenStudy (mathmale):

Let's change the approach. Suppose that there are x eligible voters, and that ALL of them vote. Shelley gets one tenth of the votes, or \[\frac{ x }{ 10 }~votes.\]

OpenStudy (mathmale):

does that make sense to you?

OpenStudy (anonymous):

we are talking about voters

OpenStudy (mathmale):

True, but we're also talking about how many votes each candidate receives. Let me try this: Suppose I tell you that there are 100 eligible voters. Suppose that all of them vote. Shelley gets one-tenth of the votes. How many votes does he get?

OpenStudy (anonymous):

we know how many votes

OpenStudy (mathmale):

Yes, we do, because we know that (1/10)th of the 100 voters voted for Shelley: thus, he got one-tenth, or 10, of the total 100 votes cast. \[\frac{ 1 }{ 10 }100=10\]

OpenStudy (anonymous):

1/2 voters in all

OpenStudy (mathmale):

Again, suppose there are 100 eligible voters and that all of them vote. Daley gets two-eighths of these votes. How many votes does Daley get? (Multipy 100 by two-eighths.)

OpenStudy (anonymous):

25

OpenStudy (mathmale):

that's right! So, if there are 100 eligible voters and all of them vote, Daley gets 25 votes and Shelley gets (1/10)(100), or 10, votes. OK with that?

OpenStudy (anonymous):

shelley 10 daley 25

OpenStudy (mathmale):

that's right! Now, suppose that only half of the 100 eligible voters actually vote. Then Shelley would receive only 1/2 of 10 votes, or 5 votes. Daley would have rec'd half of 25, or about 13 votes. OK?